OpenStudy (anonymous):

find the derivative of the function: y=(x-1)*(squareroot of 2x-x^2)+arcsin(x-1)

OpenStudy (anonymous):

you can do product rule for the first two terms and the arc sine is easy

OpenStudy (anonymous):

i dont know what to do nxt after that..

OpenStudy (anonymous):

\[y=x-1\sqrt{2x-x^2}+\arcsin(x-1)\] is this the function?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\left( x- 1\ \right)\left( 2-2x \right)\div2\sqrt{2x-x ^{2}}+\sqrt{2x-x ^{2}}+1\div \sqrt{2x-x ^{2}}\]

OpenStudy (anonymous):

is this right?